The College Board provided comparisons of Scholastic Aptitude Test (SAT) scores based on the highest level of education attained by the test taker's parents. A research hypothesis was that students whose parents had attained a higher level of education would on average score higher on the SAT. The overall mean SAT math score was 514. SAT math scores for independent samples of students follow. The first sample shows the SAT math test scores for students whose parents are college graduates with a bachelor's degree. The second sample shows the SAT math test scores for students whose parents are high school graduates but do not have a college degree.

College Grads
485 471
534 533
634 542
570 410
534 515
572 594
513 432
576 469
High School Grads
442 492
580 478
479 425
486 485
528 390
524 535

(a)

Formulate the hypotheses that can be used to determine whether the sample data support the hypothesis that students show a higher population mean math score on the SAT if their parents attained a higher level of education. (Let μ1 = population mean verbal score of students whose parents are college graduates with a bachelor's degree and μ2 = population mean verbal score of students whose parents are high school graduates but do not have a college degree.)

H0: μ1μ2 ≥ 0

Ha: μ1μ2 < 0

H0: μ1μ2 = 0

Ha: μ1μ2 ≠ 0

    

H0: μ1μ2 ≤ 0

Ha: μ1μ2 > 0

H0: μ1μ2 ≠ 0

Ha: μ1μ2 = 0

H0: μ1μ2 < 0

Ha: μ1μ2 = 0

(b)

What is the point estimate of the difference between the means for the two populations?

(c)

Find the value of the test statistic. (Round your answer to three decimal places.)

Compute the p-value for the hypothesis test. (Round your answer to four decimal places.)

p-value =

(d)

At

α = 0.05,

what is your conclusion?

Reject H0. There is sufficient evidence to conclude that higher population mean verbal scores are associated with students whose parents are college graduates.

Reject H0. There is insufficient evidence to conclude that higher population mean verbal scores are associated with students whose parents are college graduates.  

Do not Reject H0. There is sufficient evidence to conclude that higher population mean verbal scores are associated with students whose parents are college graduates.

Do not reject H0. There is insufficient evidence to conclude that higher population mean verbal scores are associated with students whose parents are college graduates.
Answer #1

H0: μ1μ2 ≤ 0

Ha: μ1μ2 > 0

b)

Point estimate =x1-x2= 37.000

c)

standard error se=√(S21/n1+S22/n2)= 21.145
test stat t =(x1-x2o)/Se= 1.750

p value =0.0462

d)

since p value <0.05:

Reject H0. There is sufficient evidence to conclude that higher population mean verbal scores are associated with students whose parents are college graduates